Erica received 14 packages. Speedy Delivery delivered twice as many as Ralph's Express, while Ralph's Express delivered two more than SendQuick Package Service. How many packages did Erica receive from each service?
Answer by rfer(16322)

sq=x=2
r=x+2=4
sd=2(x+2)=8
-------------
x+x+2+2x+4=14
4x=8
x=2